OpenStudy (anonymous):

What polynomial has roots of -4, 1, and 6?

OpenStudy (anonymous):

x3 - 3x2 - 22x + 24 x3 - x2 - 26x - 24 x3 + x2 - 26x + 24 x3 + 3x2 + 14x - 24

OpenStudy (anonymous):

cna someone help me out?

OpenStudy (anonymous):

Hey! :D

OpenStudy (anonymous):

Hey :)

OpenStudy (anonymous):

Okay so for example: (x-4)(x+5)=0 has roots x=4, -5

OpenStudy (anonymous):

You set x-4=0 and solve for x you see

OpenStudy (anonymous):

So if you know those are the roots, how can you apply this information to solve your question?

OpenStudy (anonymous):

lemme think hold on a sec

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

Hint: what are the roots of (x+4)(x-1) ?

OpenStudy (jdoe0001):

\(\large { x=\{-4,1,6\}\implies \begin{cases} x=-4\to x+4=0\to &(x+4)=0\\ x=1\to x-1=0\to &(x-1)=0\\ x=6\to x-6=0\to &(x-6)=0 \end{cases} \\ \quad \\ (x+4)(x-1)(x-6)=\textit{original polynomial} }\)

OpenStudy (anonymous):

the roots of that are 4,-1?

OpenStudy (anonymous):

would it be a?

OpenStudy (jdoe0001):

hmm dunno... multiply them all three, see what you get

OpenStudy (anonymous):

I get 24

OpenStudy (jdoe0001):

24? and the "x" went MIA then

OpenStudy (anonymous):

sorry is it x=24?

OpenStudy (jdoe0001):

ahemm.. you're meant to multiply the binomials then those will give you the original polynomial thus multiply away :)